How many ounces of 50% alcohol solution must be mixed with 60 ounces of a 20% alcohol solution to make a 30% alcohol solution?
i suck at math problems what i came up with was .50 + .20x + 60 = .30
I drew two pics on bottle with 50% but unknown ounces second bottle 20% with 60 ounces and third 30% but unknown ounces
also put X= unknown ounces Im so lost in how to put it together

Answer by mananth(16946)

-------------- percent ---------------- Amount
Alcohol I 20 ----------------60 ounces
Alcohol II 50 ----------------x ounces
Mixture 30 ---------------- 60 + x ounces

20*60+50x=30(60 +x)
1200+50x=1800+30x
50x-30x =-1200+1800
20x=600
/ 20
x = 30 ounces Alcohol II

m.ananth@hotmail.ca